Two square metal plates with sides 6.24 mm that are 5.97 mm apart have 2.17 nC of charge placed on them. What is the potential difference between the plates?
Answer #1

It is a parallel plate capacitor

A = side*side
= (6.25*10^-3)*(6.25*10^-3)
= 3.91*10^-5 m^2

d = 5.97*10^-3 m

C = ebsolene0*A/d
= (8.854*10^-12)*(3.91*10^-5)/ (5.97*10^-3)
=5.79*10^-14 F

use:
V = Q/C
= (2.17*10^-9)/(5.79*10^-14)
= 37456 V

Answer: 37456 V